OpenStudy (anonymous):

how would I do 5-2 over 0

OpenStudy (mertsj):

First you would write:

OpenStudy (mertsj):

\[\frac{5-2}{0}=\frac{3}{0}=\]

OpenStudy (mertsj):

And then, since division is the inverse of multiplication, you would figure out the answer to the corresponding multiplication example which is:

OpenStudy (mertsj):

\[0(?)=3\]

OpenStudy (mertsj):

When you figure that out, you will have the answer to your problem.

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

What is the answer?